Analysis of an accident scene indicates a car was traveling at a velocity of 67.0 mph (29.9 m/s) along the positive x-axis at the instant the brakes were applied and then traveled a distance of 115 m before hitting an obstacle at a velocity of 4.70 m/s. Determine the car's acceleration a (assumed constant) and the time t between applying the brakes and hitting the obstacle. (Assume the car is moving in the positive direction. Indicate the direction with the sign of your answer.)
Answer #1

V2 - u2 = 2 * a * S

here

S = 115 m

u = 29.9 m/s

V = 4.70 m/s

acceleration

4.70^2 - 29.9^2 = 2*a*115

a = - 3.79 m/s^2

for time

V = u+at

(V-U)/a = t

t = (4.70 - 29.9 )/(-3.79)

t = 6.64 second